OpenStudy (anonymous):

PLEASE HELP! :J What is the area of a sector with radius 10" and measure of arc equal to 45°?

OpenStudy (anonymous):

\[A.) 2.5 \pi sq. inches\]

OpenStudy (anonymous):

\[B.) 12.5 \pi sq. inches\]

OpenStudy (anonymous):

\[C.)25 \pi sq. inches\]

OpenStudy (ash2326):

@michelle_DrNiah What's the area of the circle with radius 10? leave the answer in terms of pi

OpenStudy (anonymous):

Honestly, I'm not sure :p Could you please walk me through it

OpenStudy (ash2326):

\[Area\ of\ circle = \pi\times r^2\] r= radius of the circle, can you tell the area now?

OpenStudy (anonymous):

\[Area of circle= \pi \times 100= 314\]

OpenStudy (ash2326):

leave in terms of pi so Area=100 pi

OpenStudy (ash2326):

Now this is area of circle corresponding to a sector of 360 degrees What would be the area of a sector with 45 degrees?

OpenStudy (anonymous):

12.5 pi sq. inches

OpenStudy (ash2326):

yes, good work :D did you understand?

OpenStudy (anonymous):

YAY!!! I think I do; but If I need some more help will you help you me please?

OpenStudy (ash2326):

I would if I'm on :D

OpenStudy (anonymous):

Thank you sooo much!!! :j